Fermi National Laboratory
 

Operation Summary for February 17, 2010 - February 19, 2010

 

Wednesday Feb. 17

The day shift began with Operations monitoring the Tevatron (TeV) store 7613 (76.6E30) and the Recycler stash (352.4E10), with the Antiproton Source stacking (16.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Store 7613 was terminated at 10:13 AM, operations established store 7614 at 1 PM with an Average Initial Luminosity (AIL) of 301.0E30. Shot setup ran longer than usual due to between store repairs of a TeV RF station (TRF01), some console sequencer issues, and TeV sector A3 wet engine belt tightening. Operations established beam to a new MTest experiment ( T995) at 1:43 PM. Operators reported Linac 400 MeV area Klystron Debuncher LCW flow trips.

During the evening shift, Operations called in water personnel to investigate problems with the Klystron Debuncher water skid. Operations disabled the MTest beam at 6:02 PM. The MI beam began aborting every other pulse at 7:13 PM due to a problem with the AP2 transfer line BPMs. Operators tried rebooting several devices with no luck. They contacted a Pbar expert who helped them work through a procedure that brought the BPMs back online at 8:35 PM.

Thursday Feb. 18

The midnight shift began with Operations monitoring store 7614 (91.42E30) and the stash (305.2E10), with Pbar stacking (24.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7614 at 4:07 AM and then put the TeV into shot setup. Operations established store 7615 at 5:30 AM with an AIL of 317.5E30. Operations established beam to MTest at 5:46 AM.

During the day shift, store 7615 collided all shift. PBars were stacked and transferred to the Recycler as appropriate. There was some downtime due to a new version of D59 (alarms page) causing problems. Beam was delivered to M-Test, NuMI and MiniBoone when it was available. The evening shift started with Klystron Debuncher water skid repairs. Beam resumed at 4:40 PM. Operations terminated store 7615 at 6:14 PM, CDF and D0 then made controlled accesses to fix problems. Tev A-E was accessed for PDQ cabling work. All accesses were finished by 10:44 PM, with Tev turn-on following. PBars were stacked and transferred to the Recycler as appropriate. Klystron Debuncher water skid problems caused more downtime near the end of the shift. MTest took beam until 18:00. Beam sent to NuMI and MiniBoone when it was available.

Friday Feb. 19

The midnight shift began with Operations monitoring the stash (405.5E10), with Pbar stacking (40.8E10), and with NuMI and MiniBooNE taking beam.

The Tev was in a dry squeeze. Klystron Debuncher water skid problems were traced to the controls chassis. The system was put in manual mode until the chassis could be repaired. Tev wet squeezes and orbit checks/corrections followed the dry squeeze. Shot setup was underway by 3:56 AM and store 7619 was colliding by 5:21 AM with an average initial luminosity of 306E30. MTest started taking beam at 5:24 AM. PBars were stacked and transferred to the Recycler as appropriate. Beam to NuMI and MiniBoone when it was available.

The Plans for Friday and the Weekend

The plans for today are to stack, stash and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/19/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory